# Stabilizer Entropy and entanglement complexity in the Sachdev-Ye-Kitaev model

### v3

@@ I Introduction
To further probe this transition, we analyze the entanglement spectrum and find that its adherence to Wigner-Dyson statistics is extremely sensitive to the interpolation, breaking down for both ground and high-energy states. To quantify this fragility, we employ tools inspired by quantum information theory, including the Kullback-Leibler (KL) divergence and a novel similarity measure we introduce, termed KL fidelity. These insights not only complement but also extend the findings of Ref. [14], which were based on Green’s function analysis. More specifically, we find that the critical interpolation parameter exhibits a power-law scaling, $g_{c}\sim N^{-3/4}$, and therefore vanishes in the thermodynamic limit. Thus, as $N\to\infty$, this shows quantitatively how the SYK-4 character of the ground state becomes unstable for arbitrarily small quadratic perturbations. Finally, we show that the behavior of SRE in the eigenstates can distinguish among the 888-fold symmetry classes of SYK-4. This is remarkable, as such classification is typically associated with properties of the full Hamiltonian and its eigenspectrum, rather than individual eigenstates.
@@
Figure 1: Schematic representation of the interpolated SYK-4 + SYK-2 model defined in Eq. (5). At $g=0$, the model reduces to SYK-4, characterized in the ground state by Wigner-Dyson entanglement spectrum statistics (ESS) and a near-universal (Haar-like) average capacity of entanglement, $\overline{C_{E}}$. At $g=1$, the SYK-2 limit exhibits Poisson ESS and non-universal $\overline{C_{E}}$. In the ground state, SYK-4 features are fragile under interpolation, whereas SYK-2 features remain robust. In contrast, SYK-4 behavior persists at high energies for all $g\neq 1$, rendering SYK-2 traits fragile. The capacity of entanglement serves as a probe of non-local nonstabilizerness in the eigenstates of the Hamiltonian.
@@ III.1 Entanglement
[…]
@@ III.2 Entanglement Spectrum
[…]
@@ III.4 Stabilizer Rényi Entropy
$\mathcal{M}_{\alpha}(\Psi)=\dfrac{1}{1-\alpha}\log_{2}{\left(d^{-1}\sum_{P\in\mathcal{P}_{N}}|{\rm Tr}(\Psi P)|^{2\alpha}\right)},$ (17)
the unique computable measure of non-stabilizerness [71]. The SRE can be efficiently evaluated without any minimization procedure. In Eq.17, $\Psi$ is the density operator of a $N$-qubit state $|\Psi\rangle$, $d=2^{N}$ is the Hilbert space dimension, and $\mathcal{P}_{N}$ is the $N$-qubit Pauli group. The SREs for $\alpha\geq 2$ are good resource monotones [71]. Stabilizer entropies measure how far a state deviates from stabilizer states by analyzing its spread in the Pauli operator basis [30, 72, 73].
[…]
@@
to leading order in the system size [30, 80, 64]. The factor 222 in the denominator of the linear term arises because $N/2$ represents the support of the spin/qubit representation. Our results in Eq. (18) and Eq. (19), as compared to Eq. (20), indicate that both the GS and MS of the SYK-4 model exhibit deviations from the behavior expected of fully quantum chaotic and universal states. Specifically, we observe a linear prefactor difference of approximately 0.050.050.05 relative to Haar-random states. These observations support the conclusion that the eigenstates of the interpolated SYK-4 + SYK-2 model do not fully achieve universality.
@@ Appendix D Analytical derivation of the logarithmic anti-flatness for the SYK-2 model
×$\times\ln{\left[p_{i}^{3}+(1-p_{i})^{3}\right]}$ (33)
@@
×$\times\ln{\left[\left(\dfrac{\lambda+1}{2}\right)^{3}+\left(1-\frac{\lambda+1}{2}\right)^{3}\right]}{\rm d}\lambda,$ (34)
@@
×$\times\dfrac{\sqrt{(1-\lambda^{2})+4f(1-f)-1}}{1-\lambda^{2}}{\rm d}\lambda.$ (36)
The integrable can be evaluate exactly to yield $-2(f-1)f\pi\,_{2}F_{1}\left(\frac{1}{2},1-n,2,-4(f-1)f\right)$. Using the following definition of the logarithmic anti-flatness
@@
×$\times_{2}F_{1}\left(\frac{1}{2},1-n\ ,2,4f(1-f)\right),$ (38)
@@ Appendix E Hamiltonian spectral statistics
Figure 10: Density of states (DOS) of the $H_{g}$ spectrum for $N=22$ fermions with support on 111111 qubits, with $M=40$ realizations of the disordered Hamiltonian. We considered 409604096040960 eigenvalues for each panel and corresponding $g$. See text for details and discussions. An interesting observation regarding the DOS support is highlighted below.
@@
We highlight a noteworthy aspect of the smooth change between the SYK4 and SYK2 DOS that can not be directly observed in Fig. 10. More specifically, for small values of $g$ in the range $g\in[0,0.1]$, we observe the shrinking of the energy support of the Hamiltonian eigenvalues and correspondingly the DOS support. We make this observation for all system sizes up to $N=22$ for which we computed the full spectrum. For values of $g$ that are larger than 0.10.10.1, the support grows and the DOS starts spreading rather than shrinking. This change in behavior in the DOS could been related to our observations presented in Ref. (3) for the ground state. Interestingly, the ground state properties are sensitive to these changes in the Hamiltonian behavior, while the middle of the spectrum states are not. This phenomenology can potentially be related to the findings in [14] regarding the low-temperature behavior of the interpolated SYK model.
@@
Figure 11: Spectral statistics results. Left panel: KL divergence (defined in the main text) between known Gaussian ensemble RMT distributions and the underlying dataset for the consecutive gap ratios PDF $P(r)$. Right panel: The averaged consecutive gap ratios. We considered 409604096040960 eigenvalues for each panel and corresponding $g$. Smaller systems size exhibit similar findings to the presented results for $N=22$ in this figure.
@@
Moreover, we observe that the full Hamiltonian spectrum manifestly conforms to the universal correlated statistics of RMT until around $g\leavevmode\nobreak\ \approx 0.75$ where more similarity can be found with the uncorrelated Poisson distribution (left panel). However, we note that this type of analysis is influenced by the binning of the PDF and the finite-size effects that naturally plague any finite quantum many-body system. Overall this stops us from providing conclusive judgment on the chaoticity/integrability present in the model from the full Hamiltonian spectrum. In a way, this justifies our approach to studying state complexity instead as has been done in the main text.
@@ Appendix F Analysis of the gap
In Fig. 12, we show the behavior of the average ground state to the first excited state gap for different choices of the interpolation parameter $g$. Fitting a power law function to the data we observe consistency with the expected $N^{-1}$ power law with exponent 1-1- 1 as $g$ approaches 1. On the other hand, when we approach the SYK4 point, i.e. $g$ going to 0, we are still able to fit a power law function with reasonable precision. This implies that for the finite system data, it is hard to differentiate between a power-law and the exponential behavior expected for the SYK4 model, indicating the ineffectiveness of this particular probe.
@@
Figure 14: Haar vs middle of the spectrum (MS) SYK states comparative analysis across a single bipartition (same results for other bipartitions). Upper row: Haar states included by influenced by the odd/een effects (the reason for the mismatch with the purple M-P distribution [41]); Middle row: MS state for the $g=0$ SYK4 Hamiltonian; Lower row: MS state for $g=0.01$. Other parameters: number of realization $M=100$, nb of spins $N=11$ (fermions $N=22$) and across bipartition [1,2,3,4,5][1,2,3,4,5][ 1 , 2 , 3 , 4 , 5 ].
In this section, we complete our survey of the ESS of the middle spectrum state not presented in the main text. More specifically, we repeat the same analysis presented in Fig. (4), but for the middle of the spectrum state (MS) at $E\leavevmode\nobreak\ \sim 0$, see Fig. 10. In Fig. 13, we observe that for the considered system sizes the ESS does not show agreement with RMT universal ensembles and shows good agreement with the uncorrelated Poisson distribution.
@@ Appendix G Entanglement Spectral Statistics (ESS) and normalized Reduced Density Matrix (RDM) spectrum
Appendix H Stabilizer Rényi Entropy of Middle of the spectrum states

### v4

@@ I Introduction
To further probe this transition, we analyze the entanglement spectrum and find that its adherence to Wigner-Dyson statistics is extremely sensitive to the interpolation, breaking down for both ground and high-energy states. To quantify this fragility, we employ tools inspired by quantum information theory, including the Kullback-Leibler (KL) divergence and a novel similarity measure we introduce, termed KL fidelity. These insights not only complement but also extend the findings of Ref. [14], which were based on Green’s function analysis. More specifically, we find that the critical interpolation parameter exhibits a power-law scaling, $g_{c}\sim N^{-3/4}$, and therefore vanishes in the thermodynamic limit. Thus, as $N\to\infty$, this shows quantitatively how the SYK-4 character of the ground state becomes unstable for arbitrarily small quadratic perturbations. Finally, we show that the behavior of SRE in the eigenstates can distinguish among the $8$-fold symmetry classes of SYK-4. This is remarkable, as such classification is typically associated with properties of the full Hamiltonian and its eigenspectrum, rather than individual eigenstates.
@@
Figure 1: Schematic representation of the interpolated SYK-4 + SYK-2 model defined in Eq. (5). At $g=0$, the model reduces to SYK-4, a maximally chaotic system characterized by universal entanglement properties. At $g=1$, the model reaches the SYK-2 limit which behaves like a disorder free-fermion system. In the ground state, SYK-4 features are fragile under finite values of the interpolation parameter $g$, whereas those of SYK-2 remain robust. By contrast, SYK-4 behavior persists at high energies for all $g\neq 1$, while making SYK-2 regime fragile.
@@ II The model
From the point of view of experimental realizations, synthetic quantum matter offers the most promising route to realize interpolated and other SYK-type Hamiltonians, with both analog and digital strategies actively explored in the literature. Analog proposals range from solid-state mesoscopic systems [60, 61, 62, 63, 64, 65] to cold atoms in optical lattices [66, 67, 68, 69, 70]. In particular, the authors of [64] were able to controllably engineer and enhance the SYK interactions in a disordered graphene flake using experimentally achievable magnetic fields. They achieve this result by simulating the effects of two sources of disorder: irregularities of system boundaries and bulk vacancies. On the other hand, some digital approaches based on gate-model quantum computation have also been proposed [71, 72, 73, 74]. Together, these works mark significant progress toward realizing the SYK model in quantum simulators, even if the list is not exhaustive.
@@ III.1 Entanglement
We note that due to the all-to-all connected nature of the underlying graph structure of the SYK-$q$ modes, the number of possible bipartitions grows factorially with system size, making it computationally infeasible to consider all of them when computing averages. However, the system is on average permutationally invariant so we do not need to consider all the possible bipartitions. To make the calculation numerically tractable, we instead sample $N$ bipartitions from the total of $\binom{N/2}{N/4}$ possible ones.
@@ III.2 Entanglement Spectrum
The results in the top row of Fig. 3 show the KL fidelity $D_{KL}(\eta_{g}|\eta_{g+\epsilon})$ between two distributions of the eigenvalues of the RDM for two nearby values $g,g+\epsilon$ of the interpolation parameter. This quantity can serve as a probe of a sharp transition associated with an observable consisting of a discrete probability distribution. We see that the GS and MS behave in a symmetric but opposite way. The structure of the eigenvalues of the RDM of the GS shows a sharp transition at $g>0$ and then smooths out. On the other hand, for the MS, $g=1$ is fragile, and the eigenvalues of the RDM of highly excited states are smoothly varying as long as SYK-4 interactions are different from zero. The second row of Fig. 3 shows the same statistical distance between the state for the value $g$ and the reference Haar value. While SYK-4 is converging to the Haar value, SYK-2 shoots away. The robustness of the two phases for GS and MS, respectively, is confirmed. In other words, for every value of $g\neq 1$, the eigenvalues of the RDM of highly excited states behave very similarly to the universal Haar states. The inset in Fig. 3 shows the finite-size behavior of the peaks in the KL fidelity of the MS state from the numerical data. We find that the data are well described by an exponential of the form $f(N)=1-ae^{-bN}$ with $a=0.49\,\pm\,0.08$ and $b=0.084\,\pm\,0.009$ confirming the expected value of $g_{c}=1$ [15].
@@
Experimentally, measuring the KL-fidelity requires knowledge of the reduced density matrix (RDM) eigenvalues for different values of the interpolation strength $g$. In both analog and digital simulators, this amounts to extracting the RDM spectrum of a state for a chosen bipartition, which can be done via subsystem state tomography (measuring all Pauli string correlators) [81], entanglement spectroscopy (inferring eigenvalues from RDM moments) [82], or classical shadows and randomized measurements (reconstructing the RDM from local random measurements) [83, 84]. Among these, shadow-based and randomized protocols are particularly promising for scalability, since the measurement cost grows exponentially only with the subsystem size, and remains largely independent of the full system size.
@@ III.4 Stabilizer Rényi Entropy
$\mathcal{M}_{\alpha}(\Psi)=\dfrac{1}{1-\alpha}\log_{2}{\left(d^{-1}\sum_{P\in\mathcal{P}_{N}}|{\rm Tr}(P\Psi)|^{2\alpha}\right)},$ (17)
the unique computable measure of non-stabilizerness [90]. The SRE can be efficiently evaluated without any minimization procedure. In Eq. 17, $\Psi$ is the density operator of a $N$-qubit state $|\Psi\rangle$, $d=2^{N}$ is the Hilbert space dimension, and $\mathcal{P}_{N}$ is the $N$-qubit Pauli group. The SREs for $\alpha\geq 2$ are good resource monotones [90]. Stabilizer entropies measure how far a state deviates from stabilizer states by analyzing its spread in the Pauli operator basis [30, 91, 92]. Experimental protocols to measure this quantity have recently been proposed [93, 94], allowing high-accuracy extraction of the SRE in digital quantum simulators.
In this work, we adopt the methodology outlined in Ref. [79] to compute the SRE. Specifically, we transform the GS vectors obtained from exact diagonalization into matrix product state (MPS) tensor representations and employ the Perfect Sampling algorithm [95, 96, 97]. In Fig. 6, we show the numerical results for the scaling of SRE in the GS of the $H_{g}$ model. From panel $(a)$, we observe that the SYK-4 ($g=0$) case exhibits a higher degree of non-stabilizer resources compared to the SYK-2 ($g=1$) limit. From Fig. 6 ($a$), we observe that the SRE of the SYK-4 GS does not exhibit a consistent monotonic increase with fermion number $N$. A more detailed analysis in Fig. 6 ($b$) reveals an oscillatory pattern in the SRE. The data is grouped and labeled according to the values of $N\!\!\mod 8$, as indicated in the captions. Each of the three groups corresponds to a distinct exponent of the damped exponential. This grouping is motivated by the observation in [52], where it was first noted that the SYK-4 Hamiltonian exhibits a particular particle-hole symmetry. The authors linked this symmetry to different Gaussian random matrix universality classes: GOE for $N\!\!\mod 8=0$; GUE for $N\!\!\mod 8=2,6$; and GSE for $N\!\!\mod 8=4$, which manifest in the Hamiltonian’s spectrum. We observe the effects of this symmetry specifically in the GS, without considering the full Hamiltonian spectrum. We verified numerically that this symmetry holds also in the GS. In contrast to entanglement (see Fig. 2) and ground state energy [98], the SRE is a probe of quantum complexity sensible to the finiteness of $N$. This highlights the unique ability of SRE to reveal hidden structures within many-body systems. A recent, striking example is the ability of SRE to detect a quantum phase transition that is not captured by entanglement in a system without a conventional order parameter [99].
@@
to leading order in the system size [30, 101, 79]. The factor $2$ in the denominator of the linear term arises because $N/2$ represents the support of the spin/qubit representation. Our results in Eq. (18) and Eq. (19), as compared to Eq. (20), indicate that both the GS and MS of the SYK-4 model exhibit deviations from the behavior expected of fully quantum chaotic and universal states. Specifically, we observe a linear prefactor difference of approximately $0.05$ relative to Haar-random states. These observations support the conclusion that the eigenstates of the interpolated SYK-4 + SYK-2 model do not fully achieve universality.
@@ Appendix D Analytical derivation of the logarithmic anti-flatness for the SYK-2 model
$\times\ln{\left[p_{i}^{3}+(1-p_{i})^{3}\right]}$ (33)
@@
$\times\ln{\left[\left(\dfrac{\lambda+1}{2}\right)^{3}+\left(1-\frac{\lambda+1}{2}\right)^{3}\right]}{\rm d}\lambda,$ (34)
@@
$\times\dfrac{\sqrt{(1-\lambda^{2})+4f(1-f)-1}}{1-\lambda^{2}}{\rm d}\lambda.$ (36)
The integrable can be evaluate exactly to yield $2(f-1)f\pi\,_{2}F_{1}\left(\frac{1}{2},1-n,2,-4(f-1)f\right)$. Using the following definition of the logarithmic anti-flatness
@@
$\times_{2}F_{1}\left(\frac{1}{2},1-n\ ,2,4f(1-f)\right),$ (38)
@@ Appendix E Hamiltonian spectral statistics
Figure 10: Density of states (DOS) of the $H_{g}$ spectrum for $N=22$ fermions with support on $11$ qubits, with $M=40$ realizations of the disordered Hamiltonian. We considered $40960$ eigenvalues for each panel and corresponding $g$. See text for details and discussions. An interesting observation regarding the DOS support is highlighted below.
@@
We highlight a noteworthy aspect of the smooth change between the SYK4 and SYK2 DOS that can not be directly observed in Fig. 10. More specifically, for small values of $g$ in the range $g\in[0,0.1]$, we observe the shrinking of the energy support of the Hamiltonian eigenvalues and correspondingly the DOS support. We make this observation for all system sizes up to $N=22$ for which we computed the full spectrum. For values of $g$ that are larger than $0.1$, the support grows and the DOS starts spreading rather than shrinking. This change in behavior in the DOS could been related to our observations presented in Figs. 3 and 4 for the ground state. Interestingly, the ground state properties are sensitive to these changes in the Hamiltonian behavior, while the middle of the spectrum states are not. This phenomenology can potentially be related to the findings in [14] regarding the low-temperature behavior of the interpolated SYK model.
@@
Figure 11: Spectral statistics results. Left panel: KL divergence (defined in the main text) between known Gaussian ensemble RMT distributions and the underlying dataset for the consecutive gap ratios PDF $P(r)$. Right panel: The averaged consecutive gap ratios. We considered $40960$ eigenvalues for each panel and corresponding $g$. Smaller systems size exhibit similar findings to the presented results for $N=22$ in this figure.
@@
Moreover, we observe that the full Hamiltonian spectrum manifestly conforms to the universal correlated statistics of RMT until around $g\penalty 10000\ \approx 0.75$ where more similarity can be found with the uncorrelated Poisson distribution (left panel). However, we note that this type of analysis is influenced by the binning of the PDF and the finite-size effects that naturally plague any finite quantum many-body system. Overall this stops us from providing conclusive judgment on the chaoticity/integrability present in the model from the full Hamiltonian spectrum. In a way, this justifies our approach to studying state complexity instead as has been done in the main text.
@@ Appendix F Analysis of the gap
In Fig. 12, we show the behavior of the average ground state to the first excited state gap for different choices of the interpolation parameter $g$. Fitting a power law function to the data we observe consistency with the expected $N^{-1}$ power law with exponent 1-1 as $g$ approaches 1. On the other hand, when we approach the SYK4 point, i.e. $g$ going to 0, we are still able to fit a power law function with reasonable precision. This implies that for the finite system data, it is hard to differentiate between a power-law and the exponential behavior expected for the SYK4 model, indicating the ineffectiveness of this particular probe.
@@ Appendix G Entanglement Spectral Statistics (ESS) and normalized Reduced Density Matrix (RDM) spectrum
Figure 14: Haar vs middle of the spectrum (MS) SYK states comparative analysis across a single bipartition (same results for other bipartitions). Upper row: Haar states included by influenced by the odd/een effects (the reason for the mismatch with the purple M-P distribution [41]); Middle row: MS state for the $g=0$ SYK4 Hamiltonian; Lower row: MS state for $g=0.01$. Other parameters: number of realization $M=100$, nb of spins $N=11$ (fermions $N=22$) and across bipartition $[1,2,3,4,5]$.
In this section, we complete our survey of the ESS of the middle spectrum state not presented in the main text. More specifically, we repeat the same analysis presented in Fig. (4), but for the middle of the spectrum state (MS) at $E\penalty 10000\ \sim 0$, see Fig. 10. In Fig. 13, we observe that for the considered system sizes the ESS does not show agreement with RMT universal ensembles and shows good agreement with the uncorrelated Poisson distribution.
@@
Appendix H Stabilizer Rényi Entropy of Middle of the spectrum states